1.     Given the normally distributed variable X with mean 18 and variance 6.25. find

a)   P(X is less than15)

b)   THE VALUE OF K SUCH THAT P(X is less than k) = 0.2236

c)    P(17 less than X less than 21)


1
Expert's answer
2021-05-31T16:04:22-0400

"The \space random\space variable \space X \space follows \space normal\ distribution \space\\ with \space a \space mean\space \mu \space =18 \space ,variance \space =6.25 \space and \space the \space standard\space deviation \space of \space \sigma=2.5\\ 
(a)\\ 

Lets \space find \space P(X<15)\\
First, \space we \space we\space have \space to \space find\space z\space value \space corresponding \space to \space x=15\\
z=\frac{x-\mu}{\sigma}=\frac{15-18}{2.5}=-1.2\\
hence\\
P(X<15)=P(Z<-1.2)=0.1151\\
put \space value \space from\space z \space table\\
(b)\\
Lets \space find \space the \space value\space of \space k\space such\space that \space P(X)=0.2236\\
P(z<\frac{k-\mu}{\sigma})=0.2236\\
P(z<-0.76)=0.2236\\
put \space value \space from\space z \space table\\
hence\\
\frac{k-\mu}{\sigma}=-0.76\\
k=-0.76\sigma +\mu\\
k=-0.76×2.5 +18\\
k=16.1\\
(c)
\\
Lets\space find\space P(17<X<21)\\
first, \space we \space have \space to \space find\space z\space values\space corresponding \space to \space x_1=17 \space and \space x_2=21\\
z_1=\frac{x_1-\mu}{\sigma}=\frac{17-18}{2.5}=-0.4\\
z_2=\frac{x_2-\mu}{\sigma}=\frac{21-18}{2.5}=1.2\\
hence\\
P(17<X<21)=P(-0.4<Z<1.2)\\
=P(Z<1.2)-P(Z<-0.4)
put \space value \space from\space z \space table\\
=0.8849-0.3446
=0.5403"